OpenStudy (comm.dan):

Who is good with inequalities? Because I need help. I will give a medal to the person I believe answers my question I distributed first Here is the question: x(7-x)>8 I distributed to get 7x-x>8 I don't know what to do after this

OpenStudy (anonymous):

you didn't distribute... \[x(7-x)>8 \Rightarrow 7x-x^{2}>8\]

OpenStudy (comm.dan):

That is what I meant

OpenStudy (comm.dan):

@pgpilot326

OpenStudy (anonymous):

som next, get everything on one side and factor...

OpenStudy (anonymous):

so

OpenStudy (isaiah.feynman):

Okay then, this leads to a quadratic equation.

OpenStudy (comm.dan):

\[7x-x squared >8\]

OpenStudy (comm.dan):

There are going to be two solutions Here are the choices 2,5,-1, and zero

OpenStudy (anonymous):

move everything to the RHS \[0>x^{2}-7x+8\] now factor the only way the inequality can be satisfied is if the RHS is negative

OpenStudy (comm.dan):

Right side?

OpenStudy (anonymous):

oops. same idea but correct factoring.

OpenStudy (comm.dan):

How did u get the above answer?

OpenStudy (e.mccormick):

They did the right side to change the \(-x^2\) into a positive. They could have also done it on the left side by turning the > into a < and multiplying everything by \(-1\).

OpenStudy (anonymous):

sorry, it doesn't factor like that. once factored, plot the zeros on a number line. you'll get 3 regions. see if any of those regions satisfy the inequality

OpenStudy (anonymous):

you can always use the quadratic formula

OpenStudy (comm.dan):

Multiply everything by -1, the inequality sign switches directions, and then you solve the equation regularly?

OpenStudy (anonymous):

yeah or just rewrite... \[0>x^{2}-7x+8 \Rightarrow x^{2}-7x+8<0\] Come on @e.mccormick

OpenStudy (e.mccormick):

Just allows you to change the left and right and get it into a format you might be a little more comfortable with.

OpenStudy (comm.dan):

What should I do next?

OpenStudy (e.mccormick):

Do you understand how you will use the quadratic in this problem? pgpilot326 is quite right in saying you will need it.

OpenStudy (comm.dan):

Make it into parenthesis (x+#) (x+#) Something like this

OpenStudy (comm.dan):

It would be x- x- a number for each

OpenStudy (e.mccormick):

Yes. So put \(x^{2}-7x+8=0\) into the quadratic like you normally would. Get the exact answers. Then put the exact answers in where the equation was on the original: \(x^{2}-7x+8<0\implies (x\pm ?_1)(x\pm ?_2)<0 \)

OpenStudy (comm.dan):

I am thinking of x-2 x-5 and it isn't working because I remember that both of the numbers have to add up to -7x and when multiplied together equal positive 8

OpenStudy (e.mccormick):

Which is why you need the quadratic. You will have a root.

OpenStudy (anonymous):

\[\frac{-b \pm \sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (comm.dan):

Do I do square root signs for both of them?

OpenStudy (e.mccormick):

@pgpilot326 And yes, flipping it after getting it with the > works very well too! I was just pointing out the math of doing so. @Comm.Dan Plug in the numbers and see.

OpenStudy (comm.dan):

K Thx for the reminder of the formula I forgot it because I haven't done quadratics in a long time

OpenStudy (comm.dan):

@pgpilot326

OpenStudy (comm.dan):

Thx for the help so far @pgpilot326 and @e.mccormick

OpenStudy (anonymous):

@e.mccormick i understand but really there doesn't need to be any math involved. saying that 0 > -1 is the same as saying that -1 < 0. there is math to rework but i think it's better to understand the simplicity first.

OpenStudy (anonymous):

don't get me wrong, what you did is perfectly valid and I mean no disrespect @e.mccormick

OpenStudy (comm.dan):

\[7\pm \sqrt-39\] over 2

OpenStudy (e.mccormick):

-39?

OpenStudy (e.mccormick):

\((-7)^2-4(1)(8)\)

OpenStudy (comm.dan):

Sorry, my bad I looked and I made a mistake

OpenStudy (e.mccormick):

Yah, forgot to square it first. =)

OpenStudy (anonymous):

? \[\frac{7 \pm \sqrt{(-7)^{2}-4(1)(8)}}{2(1)}=\frac{7 \pm \sqrt{49-32)}}{2}=\frac{7 \pm \sqrt{17)}}{2}\]

OpenStudy (e.mccormick):

I got: \(\dfrac{7+\sqrt{17}}{2},\dfrac{7-\sqrt{17}}{2}\) You should too after fixing it.

OpenStudy (comm.dan):

I did, too. I got the same after realizing my mistake

OpenStudy (e.mccormick):

OK, so: \(\left(x-\dfrac{7+\sqrt{17}}{2}\right)\left(x-\dfrac{7-\sqrt{17}}{2}\right) > 0\) Where is that true? Try a value above, between and below.

OpenStudy (anonymous):